Soybean meal is 18% protein and cornmeal is 9% protein. How many pounds of each mixture should be mixed together in order to get 360lb mixture that is 16% protein.
Answer by helpnalgebra(91)

this has been answered here it is for you.
.18X+.09(360-X)=.16*360
.18X+32.4-.09X=57.6
.09X=57.6-32.4
.09X=25.2
X=25.2/.09
X=280 LBS. OF SOYBEANS ARE USED.
360-280=80 LBS. OF CORNMEAL ARE USED.
PROOF:
.18*280+.09*80=57.6
50.4+7.2==57.6
57.6=57.6